OpenStudy (anonymous):

15% (15/100)= .15 (.15 X 120) + 120

OpenStudy (callisto):

Hmm.. what is your question?

OpenStudy (anonymous):

this is my question : Karim borrowed $120 from his older brother for the purchase of a skateboard. He promised to pay back the loan in 6 months at a simple annual interest rate of 15%. How much will Karim owe his brother at that time? - and 15% (15/100)= .15 (.15 X 120) + 120 is what i got so far im stuck and not sure what to do next.

OpenStudy (callisto):

Interest he needs to pay back = 120 x 15% x 6/12 = 9 Amount he owes = 120 + interest = 120 +9 =?

OpenStudy (anonymous):

129$